OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5 Help please.

jhonyy9 (jhonyy9):

like a first step note sinx = t and solve this quadratic for t so when you get roots rewrite it for t = sinx than x_1 = arcsin t_1 and x_2 = arcsin t_2 hope this will help you

OpenStudy (mertsj):

Is it 7sin(2x) or 7sin^2x ???

OpenStudy (anonymous):

It is 7sin^2x

OpenStudy (anonymous):

@jhonyy9 I am sure it does, but I am not good at math whatsoever therefore your answer was a bit complex for me to understand. Thanks for replying though! It is very appreciated.

OpenStudy (mertsj):

\[7\sin ^2x-14\sin x+2=-5\] \[7\sin ^2x-14\sin x+7=0\] \[\sin ^2x-2\sin x+1=0\]

OpenStudy (mertsj):

Factor. Set each factor equal to 0 and solve.

OpenStudy (anonymous):

Thank you so much! Have a good night!

OpenStudy (mertsj):

yw